1) Consider and A.P with a as its first term amd d as its common difference.Let S k denote the sum of its first K terms. If S kx /S x is independent of x, then prove that a=d/2.

1)Consider and A.P with a as its first term amd d as its common difference.Let Sk denote the sum of its first K terms. If Skx/Sx is independent of x,  then prove that a=d/2.
 

Grade:11

1 Answers

Sourabh Singh IIT Patna
9 years ago
Hii


Sk= k/2(2*a + (k-1)*d) and


Sx= x/2(2*a + (x-1)*d)

If after dividing it we are getting something which is independent of “x” then we will get a relation betwen the first term and the common difference.

i.e. a=d/2.

Thanks